Session: Jun 2011 (session 2)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (Bo Xia at Fri May 27 04:13:50 2011) Success:

Medicina (C.Bortolotti at Fri May 27 04:32:55 2011) Minor failures:
VC01 e VC02 alarm for very few time.

Onsala (Michael Lindqvist at Fri May 27 08:59:34 2011) Success:
Raining first part of the experiment. Otherwise no known problems.

Torun (Kaz Borkowski at Fri May 27 12:30:37 2011) Minor failures:
Recording off source between 21:53 and 22:08 UTC because of problems with pumps of cryo system earlier this day (compressor unexpectantly stopped working).

Yebes (Pablo de Vicente at Fri May 27 12:59:34 2011) Success:
No known problems

Westerbork (Geert Kuper at Sat May 28 12:35:27 2011) Minor failures:
Note that WB needs at least 3 minutes to initialize its backend. This is needed after each frequency change, so some of the scans in this project may be shortened or omitted. Scans: 11,117,139,161,162,182,183,204,205,224,225,245,246,271

Effelsberg (Uwe Bach at Mon May 30 10:33:08 2011) Success: